OpenStudy (istim):

Simple dimensional analysis: To cancel out "s" in J/S, do I multiply or divide J/s by s? s in both cases are the same.

OpenStudy (istim):

This is part of a physics question, but I thought it was simple algebra, and deserved to be let free here.

OpenStudy (whpalmer4):

joules are kg m^2/s^2 \[J/s= \frac{{kg}*m^2}{s^2}*\frac{1}{s}=\frac{{kg}*m^2}{s^3}\] Right?

OpenStudy (istim):

I think this is extremely over-complicating the original intent of this thread...but...yes?

OpenStudy (istim):

Simply asking how to cancel out "s" in this situation.

OpenStudy (whpalmer4):

Where are you canceling the s? Maybe with more context I can give you a better answer.

OpenStudy (istim):

In J/S

OpenStudy (whpalmer4):

They don't cancel out. J/s = watts, and watts are kg m^2 / s^3, just like I showed above.

OpenStudy (whpalmer4):

Remember, dividing by s is equivalent to multiplying by 1/s.

OpenStudy (istim):

well, you know, the same well you can "remove" distance (m) in speed (m/s) to get time (t) by

OpenStudy (istim):

t=v/d? t=d/v?

OpenStudy (istim):

Man, can not realize I forgot this so fast.

OpenStudy (whpalmer4):

d = vt m = (m/s)*(s) m = m But there's no cancellation in J/s...